### v1

@@ 4.2 Validation
Figure 6: Comparison between the integral EW and the fitted EW of the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ for their HQ samples, from the sub-figure (a) to (c). For each sub-figure, the upper panel shows the Gaussian KDE number density plot of the integral EW as a function of the fitted EW, with the red dashed line representing the one-to-one relation. The lower panel displays the Gaussian KDE number density plot of the difference between the integral EW and the fitted EW, i.e., $\rm\Delta EW=EW_{int}-EW_{fit}$ as a function of the fitted EW, with the green dashed line representing the zero benchmark. The distribution of $\rm\Delta EW$ is also overplotted in the upper panel with its mean and standard deviation annotated.
@@ 4.2.3 Multiple Epochs
Benefiting from the high efficiency in the spectrum acquisition, LAMOST has accumulated numerous sources with multiple epochs. These sources that have been observed many times can provide a good opportunity to validate the variation of DIB measurement. We select the sources with over two epochs from the HQ samples by grouping the sources with the same uid (unique source identifier in LAMOST) and present the distribution of the number of epochs in the panels in the first column of Figure 7. There are 15,839, 891 and 9,939 sources with over two epochs for the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$, respectively. These sources have four epochs on average and the maximum number of epochs is up to 33. The panels in the second column of Figure 7 illustrate the cumulative distribution of the standard deviation of the EWs ($\rm\sigma_{EW}$) for multiple-epoch sources. Almost 95% of the sources have $\rm\sigma_{EW}<0.05$ for three DIBs, which is comparable to the typical value, i.e., 0.04 Å of $\rm\Delta EW$ investigated in Section 4.2.2. Such similarity between $\rm\sigma_{EW}$ and $\rm\Delta EW$ implies that the variation of the DIB measurements among multiple epochs of the same source is resulted from the measurement error rather than the intrinsic property of the DIB. Without the influence of blend with a broader DIB on the DIB $\lambda 6614$ (see Section 4.2.2), it is not surprising that the cumulative distribution of the DIB $\lambda 6614$ accumulates toward 1 more rapidly than that of the DIBs $\lambda 5780$ and $\lambda 5797$.
@@
In order to present the relative variation of the EWs, we also supply the cumulative distribution of the coefficient of variation (CV = $\rm\sigma_{EW}/EW_{mean}$) in the panels in the third column of Figure 7. Well over three quarters of the sources have CV $<0.2$ for the DIB $\lambda 5797$, and this proportion is more than 90% for the DIBs $\lambda 5780$ and $\lambda 6614$. The large CV of the multiple-epoch sources generally comes from the small EW, which appears to be more sensitive to the measurement error, as can be seen in the panels in the fourth column of Figure 7. It is noted that both the cumulative distribution of the $\rm\sigma_{EW}$ and the CV demonstrate the long tail. Besides the measurement error mentioned above, such an imbalance is driven by the differences among the atmospheric parameters of multiple spectra for the same source. Specifically, although these spectra originate from the same source, the differences in observation conditions, processing procedures, instrument running status, etc., result in variations in atmospheric parameter measurements among these spectra. Consequently, their matched closest neighbor template spectra also differ, which ultimately leads to a noticeable discrepancy in the ISM residual spectra. In the visual inspection, we found that the maximum difference in $T_{\rm eff}$ among multiple-epoch spectra of the same source can reach 500 K, far exceeding the match threshold of 100 K set in Section 3.2.1. Zhao et al. (2024) and Saydjari et al. (2023) have circumvented the issues of atmospheric parameters and directly extracted the DIB $\lambda 8621$ signal from the parameter-free spectra using machine learning methods. The applicability of this method, however, requires further validation, which is also one of our future directions.
@@
Figure 7: Three distinct colors encode the three DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ for the HQ samples with over two epochs. The panels in the first column present the distribution of the number of epochs. The panels in the second column depict the cumulative distribution of standard deviation of the EWs ($\rm\sigma_{EW}$). The panels in the third column display the cumulative distribution of the coefficient of variation ($\rm\sigma_{EW}/EW_{mean}$). The panels in the fourth column show the scatter of the $\rm\sigma_{EW}/EW_{mean}$ as a function of the $\rm EW_{mean}$.
@@ 4.2.4 Residuals
In addition to validating the measurements of DIBs, the quality of ISM residual spectra is also verified. Ideally, after subtracting the stellar features using the template spectrum, the flux of the ISM residual spectrum outside the DIB region which is from $\rm\mu_{mc}-3\sigma_{mc}$ to $\rm\mu_{mc}+3\sigma_{mc}$ should be at the continuum level with no stellar components. Thus, we calculate the standard deviation of the residuals between the flux of ISM residual spectrum outside the DIB region and that of the continuum, and then provide the trend of STD with SNR_r and $T_{\rm eff}$ in Figure 8. As can be seen in the figure, the STD decreases as SNR_r increases, which indicates that higher SNR_r leads to smaller residuals in the ISM residual spectrum or cleaner subtraction of stellar components. The STD also decreases with increasing $T_{\rm eff}$ as a consequence of the weakening of metal lines in hotter stars. However, this decreasing trend of $T_{\rm eff}$ is more gradual compared to that of SNR_r, reflecting the robustness of measurement, which is less disturbed by absorption lines in cool stars. Furthermore, the magnitude of STD is comparable to the level of normalized flux error estimated in 3.2.2, which indicates that the pipeline does not introduce additional errors and that the ISM residual spectrum is quite clean, with no significant residual stellar components. It is noteworthy that although there probably exist the DIB components other than the interesting ones within the range of the ISM spectrum, they are usually weak absorptions, e.g., the DIBs $\lambda 5762$ and $\lambda 5815$ in Figure 3, documented in the previous DIB catalogs (Fan et al. 2019; Vogrinčič et al. 2023). Moreover, after instrumental broadening by the LAMOST LRS, the depths of those weak DIBs and the flux errors share the close values, which cannot cause any significant bias in the final results.
Figure 8: Trend of the standard deviation of the residuals between the flux of ISM residual spectrum beyond the DIB region and that of the continuum as a function of SNR in $r$ band in the upper panels and $T_{\rm eff}$ in the lower panels. Three distinct colors encode the three DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$. In each panel, a violin plot is used to present the distribution of the standard deviation of the residuals for the binned HQ samples. The median of each bin is annotated by the white dot, and the 2.5 and 97.5 percentiles are marked by the black dots. The red dashed line connects the median of each bin to show the trend of the standard deviation of the residuals.
@@ 5 Results and Discussion
In this section, we review and explore the spatial distribution, correlation, central wavelength and kinematics of the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ in the context of statistics based on the HQ samples. Due to the multiple-epoch sources mentioned in Section 4.2.3, we combine the measurements of DIB from the same source and take the mean value as its final result. After combination, the number of the HQ sources for the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ is 142,074, 11,480 and 85,301 respectively.
@@ 5.1 Spatial Distribution
Figure 9 presents the number density of the background stars and the sky distribution of the median EW of the DIBs $\lambda 5780$ and $\lambda 6614$. The absence of the DIB $\lambda 5797$ in the figure is due to the scarcity of the HQ sources caused by the low detection rate (see Table 3). In order to present a more generalized sky distribution of DIB in the lower panels of Figure 9, we remove the pixels in which the number density of the background stars is less than 3. Among the removed pixels, some exhibit significantly larger EWs than the values of DIB in the surrounding areas. Upon inspecting the spectra of the DIBs with large EW, we find that their background stars are located at considerable distances, which can lead to the large extinctions and EWs. However, there are some large EWs of DIB at the sight-line of the short-distance background stars as well, which are likely to be the fake DIB signals investigated in Section 4 or the incomplete correlation between the DIB and extinction discussed in Section 5.2.1.
@@
Figure 9: Galactic distribution of the DIBs $\lambda 5780$ and $\lambda 6614$ for the HQ samples. The upper panels show the number density of the background stars. The lower panels display the sky distribution of the strength of the DIBs with the color bar representing the median EW in each pixel. The black contours of the corresponding median $\rm E(B-V)$ at the level of 0.35 and 0.7 magnitude are overplotted in the median EW map. The resolution of map is about 0.92 ($N_{\rm side}=64$) calculated by the HEALPix package.
@@
The DIBs $\lambda 5780$ and $\lambda 6614$ share a similar sky distribution, that is, the Galactic plane is the prominent region with the strong DIBs and the strength of the DIBs decreases with the increasing Galactic latitude. The lower number of samples for DIB $\lambda 6614$ compared to DIB $\lambda 5780$ at high latitudes is caused by the weaker intensity of DIB $\lambda 6614$ per unit extinction (Vogrinčič et al. 2023), making it less detectable in the regions of low extinction at high galactic latitudes. The high-EW regions of the DIBs $\lambda 5780$ and $\lambda 6614$, under the Galactic view, are both spatially correlated with the well-known molecular clouds, such as Taurus and Orion. This behavior is similar to that of the DIB $\lambda 8621$ investigated in $Gaia$ RVS spectra (Gaia Collaboration et al. 2023b; Saydjari et al. 2023). In addition, there is a notable extent to high latitudes ($b\approx-45^{\circ}$) in the directions of Galactic anti-center, which does not entirely align with the distribution of molecular clouds. This stretch from the plane to high latitudes is similar to that of the DIB $\lambda 8621$ depicted by Gaia Collaboration et al. (2023b), Saydjari et al. (2023) and Zhao et al. (2024). The causes of such a phenomenon remain unclear, but it is likely attributable to the selection effect of the observation, as can be seen in the sky distribution of the number density in the upper panels of Figure 9. Another possible explanation is that the resolution of map is not high enough to resolve the sky distribution of the DIBs, which makes it appear as though the region is interconnected. It is anticipated that with the inclusion of DIB measurements from hot stars of LAMOST in the future, more detailed and accurate northern sky distribution of the DIBs will be obtained. Additionally, a three-dimensional spatial distribution of the DIBs, e.g., Kos et al. (2014) and Farhang et al. (2019), will be constructed, which can provide a more comprehensive understanding of the DIBs and their positional and physical relationship with the ISM.
@@ 5.2 Correlation
There have been numerous studies on the correlation of DIB with DIB and DIB with the ISM tracers such as extinction, CO gas, atomic hydrogen and molecular hydrogen (e.g., Herbig 1975; Weselak et al. 2004; Friedman et al. 2011; Lan et al. 2015). However, these works are more or less limited by the number of sight-lines or combined the spectra along the same sight-line for high SNR, which sacrifice the statistical generalization. For the sake of statistical and homogeneous comparison, a total of 7,681 common samples for the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ from the HQ sources are picked out to investigate the correlation of DIB V.S. extinction and DIB V.S. DIB.
5.2.1 DIB V.S. E(B-V)
To avoid zero or inaccurate values of $\rm E(B-V)$ at close distances, the heliocentric distance $d$ is limited to greater than 500 pc and a total of 6,514 samples are selected. Figure 10 compares the EW of DIB with the extinction $\rm E(B-V)$ derived from Green et al. (2019) by the linear fit to the median EW in each $\rm\Delta E(B-V)=0.05\ mag$ bin. All three DIBs agree with a linear positive correlation with $\rm E(B-V)$ in Table 4, but not perfect, as known for a few decades (Herbig 1995). Nonetheless, the areas with high number density (in red) show smaller scatters and have better linear relationship (red solid lines). Even in sparsely sampled regions, such as those near low extinction of 0.15 mag and high extinction of over 0.6 mag, the medians do not deviate from the fitted curve. We do not include the samples with the higher extinction ($>$ 1 mag) in the fitting process due to the well-known “skin effect” (Herbig 1995), namely, the strength of DIB cannot indefinitely increase with the extinction as expected. Besides, we also attempt a linear fit forced through the origin (green dashed lines) to facilitate the comparison with the previous works, e.g., Herbig (1975). It is clear that the green dashed lines are not suitable to represent the correlation of the DIBs with $\rm E(B-V)$, as they have larger fitting errors and deviate from the red solid lines.
Figure 10: Gaussian KDE number density plot of the EWs of the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ as a function of the extinction $\rm E(B-V)$ derived from Green et al. (2019), from the top panel to the bottom panel. For each panel, the red solid line represents a linear fit to the median EW (marked in white dot and its standard deviation marked in black error bar) in each $\rm\Delta E(B-V)=0.05\ mag$ bin. The green dashed line is the same as the red solid line, but the intercept is forced to be zero. The coefficients of the linear fit and the Pearson correlation coefficient between the EW and $\rm E(B-V)$ are also annotated in each panel.
Table 4: Correlation of DIB with $\rm E(B-V)$ $(1)$$(1)$footnotemark: $(1)$

### v2

@@ 4.2.2 Integral EW versus fitted EW
Figure 6: Comparison between the integral EW and the fitted EW of the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ for their HQ samples, from the subfigures (a) to (c). For each subfigure, the upper panel shows the Gaussian KDE number density plot of the integral EW as a function of the fitted EW, with the red dashed line representing the one-to-one relation. The lower panel displays the Gaussian KDE number density plot of the difference between the integral EW and the fitted EW, namely $\rm\Delta EW=EW_{int}-EW_{fit}$ as a function of the fitted EW, with the green dashed line representing the zero benchmark. The distribution of $\rm\Delta EW$ is also overplotted in the upper panel with its mean and standard deviation annotated.
@@ 4.2.3 Multiple epochs
Benefiting from the high efficiency in the spectrum acquisition, LAMOST has accumulated numerous sources with multiple epochs. These sources that have been observed many times can provide a good opportunity to validate the variation of DIB measurement. We select the sources with over two epochs from the HQ samples by grouping the sources with the same uid (unique source identifier in LAMOST) and present the distribution of the number of epochs in the panels in the first column of Fig. 7. There are 15,839, 891, and 9,939 sources with over two epochs for the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$, respectively. These sources have four epochs on average and the maximum number of epochs is up to 33. The panels in the second column of Fig. 7 illustrate the cumulative distribution of the standard deviation of the EWs ($\rm\sigma_{EW}$) for multiple-epoch sources. Almost 95% of the sources have $\rm\sigma_{EW}<0.05$ for three DIBs, which is comparable to the typical value, namely 0.04 Å of $\rm\Delta EW$ investigated in Sect. 4.2.2. Such similarity between $\rm\sigma_{EW}$ and $\rm\Delta EW$ implies that the variation of the DIB measurements among multiple epochs of the same source is resulted from the measurement error rather than the intrinsic property of the DIB. Without the influence of blend with a broader DIB on the DIB $\lambda 6614$ (see Sect. 4.2.2), it is not surprising that the cumulative distribution of the DIB $\lambda 6614$ accumulates toward 1 more rapidly than that of the DIBs $\lambda 5780$ and $\lambda 5797$.
@@
In order to present the relative variation of the EWs, we also supply the cumulative distribution of the coefficient of variation (CV = $\rm\sigma_{EW}/EW_{mean}$) in the panels in the third column of Fig. 7. Well over three quarters of the sources have CV $<0.2$ for the DIB $\lambda 5797$, and this proportion is more than 90% for the DIBs $\lambda 5780$ and $\lambda 6614$. The large CV of the multiple-epoch sources generally comes from the small EW, which appears to be more sensitive to the measurement error, as can be seen in the panels in the fourth column of Fig. 7. It is noted that both the cumulative distribution of the $\rm\sigma_{EW}$ and the CV demonstrate the long tail. Besides the measurement error mentioned above, such an imbalance is driven by the differences among the atmospheric parameters of multiple spectra for the same source. Specifically, although these spectra originate from the same source, the differences in observation conditions, processing procedures, instrument running status, etc., result in variations in atmospheric parameter measurements among these spectra. Consequently, their matched closest neighboring template spectra also differ, which ultimately leads to a noticeable discrepancy in the ISM residual spectra. In the visual inspection, we found that the maximum difference in $T_{\rm eff}$ among multiple-epoch spectra of the same source can reach 500 K, far exceeding the match threshold of 100 K set in Sect. 3.2.1. Zhao et al. (2024) and Saydjari et al. (2023) have circumvented the issues of atmospheric parameters and directly extracted the DIB $\lambda 8621$ signal from the parameter-free spectra using machine learning methods. The applicability of this method, however, requires further validation, which is also one of our future directions.
@@
Figure 7: Three distinct colors encode the three DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ for the HQ samples with over two epochs. The panels in the first column present the distribution of the number of epochs. The panels in the second column depict the cumulative distribution of the standard deviation of the EWs ($\rm\sigma_{EW}$). The panels in the third column display the cumulative distribution of the coefficient of variation ($\rm\sigma_{EW}/EW_{mean}$). The panels in the fourth column show the scatter of the $\rm\sigma_{EW}/EW_{mean}$ as a function of the $\rm EW_{mean}$.
@@ 4.2.4 Residuals
In addition to validating the measurements of DIBs, the quality of ISM residual spectra is also verified. Ideally, after subtracting the stellar features using the template spectrum, the flux of the ISM residual spectrum outside the DIB region which is from $\rm\mu_{mc}-3\sigma_{mc}$ to $\rm\mu_{mc}+3\sigma_{mc}$ should be at the continuum level with no stellar components. Thus, we calculate the standard deviation of the residuals between the flux of ISM residual spectrum outside the DIB region and that of the continuum, and then provide the trend of STD with S/N and $T_{\rm eff}$ in Fig. 8. As can be seen in the figure, the STD decreases as S/N increases, which indicates that higher S/N leads to smaller residuals in the ISM residual spectrum or cleaner subtraction of stellar components. The STD also decreases with increasing $T_{\rm eff}$ as a consequence of the weakening of metal lines in hotter stars. However, this decreasing trend of $T_{\rm eff}$ is more gradual compared to that of S/N, reflecting the robustness of measurement, which is less disturbed by absorption lines in cool stars. Furthermore, the magnitude of STD is comparable to the level of normalized flux error estimated in 3.2.2, which indicates that the pipeline does not introduce additional errors and that the ISM residual spectrum is quite clean, with no significant residual stellar components. It is noteworthy that although there probably exist the DIB components other than the interesting ones within the range of the ISM spectrum, they are usually weak absorptions, such as the DIBs $\lambda 5762$ and $\lambda 5815$ in Fig. 3, documented in the previous DIB catalogs (Fan et al. 2019; Vogrinčič et al. 2023). Moreover, after instrumental broadening by the LAMOST LRS, the depths of those weak DIBs and the flux errors share the close values, which cannot cause any significant bias in the final results.
Figure 8: Trend of the standard deviation of the residuals between the flux of the ISM residual spectrum beyond the DIB region and that of the continuum as a function of S/N in the upper panels, and $T_{\rm eff}$ in the lower panels. Three distinct colors encode the three DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$. In each panel, a violin plot is used to present the distribution of the standard deviation of the residuals for the binned HQ samples. The median of each bin is annotated with a white dot, and the 2.5 and 97.5 percentiles are marked with black dots. The red dashed line connects the median of each bin to show the trend of the standard deviation of the residuals.
@@ 5 Results and discussion
In this section, we review and explore the spatial distribution, correlation, central wavelength and kinematics of the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ in the context of statistics based on the HQ samples. Due to the multiple-epoch sources mentioned in Sect. 4.2.3, we combine the measurements of DIB from the same source and take the mean value as its final result. After combination, the number of the HQ sources for the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ is 142,074, 11,480, and 85,301 respectively.
@@ 5.1 Spatial distribution
Figure 9 presents the number density of the background stars and the sky distribution of the median EW of the DIBs $\lambda 5780$ and $\lambda 6614$. The absence of the DIB $\lambda 5797$ in the figure is due to the scarcity of the HQ sources caused by the low detection rate (see Table 3). In order to present a more generalized sky distribution of DIB in the lower panels of Fig. 9, we remove the pixels in which the number density of the background stars is less than 3. Among the removed pixels, some exhibit significantly larger EWs than the values of DIB in the surrounding areas. Upon inspecting the spectra of the DIBs with large EW, we find that their background stars are located at considerable distances, which can lead to the large extinctions and EWs. However, there are some large EWs of DIB at the sight line of the short-distance background stars as well, which are likely to be the fake DIB signals investigated in Sect. 4 or the incomplete correlation between the DIB and extinction discussed in Sect. 5.2.1.
@@
Figure 9: Galactic distribution of the DIBs $\lambda 5780$ and $\lambda 6614$ for the HQ samples. The upper panels show the number density of the background stars. The lower panels display the sky distribution of the strength of the DIBs with the color bar representing the median EW in each pixel. The black contours of the corresponding median $\rm E(B-V)$ at the level of 0.35 and 0.7 magnitude are overplotted in the median EW map. The resolution of the map is about 0.92 ($N_{\rm side}=64$) calculated by the HEALPix package.
@@
The DIBs $\lambda 5780$ and $\lambda 6614$ share a similar sky distribution, that is, the Galactic plane is the prominent region with the strong DIBs and the strength of the DIBs decreases with the increasing Galactic latitude. The lower number of samples for DIB $\lambda 6614$ compared to DIB $\lambda 5780$ at high latitudes is caused by the weaker intensity of DIB $\lambda 6614$ per unit extinction (Vogrinčič et al. 2023), making it less detectable in the regions of low extinction at high galactic latitudes. The high-EW regions of the DIBs $\lambda 5780$ and $\lambda 6614$, under the Galactic view, are both spatially correlated with the well-known molecular clouds, such as Taurus and Orion. This behavior is similar to that of the DIB $\lambda 8621$ investigated in $Gaia$ RVS spectra (Gaia Collaboration et al. 2023b; Saydjari et al. 2023). In addition, there is a notable extent to high latitudes ($b\approx-45^{\circ}$) in the directions of Galactic anti-center, which does not entirely align with the distribution of molecular clouds. This stretch from the plane to high latitudes is similar to that of the DIB $\lambda 8621$ depicted by Gaia Collaboration et al. (2023b), Saydjari et al. (2023), and Zhao et al. (2024). The causes of such a phenomenon remain unclear, but it is likely attributable to the selection effect of the observation, as can be seen in the sky distribution of the number density in the upper panels of Fig. 9. Another possible explanation is that the resolution of map is not high enough to resolve the sky distribution of the DIBs, which makes it appear as though the region is interconnected. It is anticipated that with the inclusion of DIB measurements from hot stars of LAMOST in the future, more detailed and accurate northern sky distribution of the DIBs will be obtained. Additionally, a three-dimensional spatial distribution of the DIBs, such as Kos et al. (2014) and Farhang et al. (2019), will be constructed, which can provide a more comprehensive understanding of the DIBs and their positional and physical relationship with the ISM.
@@ 5.2 Correlation
There have been numerous studies on the correlation of DIB with DIB and DIB with other ISM tracers, such as extinction, CO gas, atomic hydrogen, and molecular hydrogen (e.g., Herbig 1975; Weselak et al. 2004; Friedman et al. 2011; Lan et al. 2015). However, these works are more or less limited by the number of sight lines or combined the spectra along the same sight line for high S/N, which sacrifice the statistical generalization. For the sake of statistical and homogeneous comparison, a total of 7,681 common samples for the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ from the HQ samples are picked out to investigate the correlation of DIB V.S. extinction and DIB V.S. DIB.
5.2.1 Correlation between DIBs and E(B-V)
To avoid zero or inaccurate values of $\rm E(B-V)$ at close distances, the heliocentric distance $d$ is limited to greater than 500 pc and a total of 6,514 common HQ sources are selected. Figure 10 compares the EW of DIB with the extinction $\rm E(B-V)$ derived from Green et al. (2019) by the linear fit to the median EW in each $\rm\Delta E(B-V)=0.05\ mag$ bin. All three DIBs agree with a linear positive correlation with $\rm E(B-V)$ in Table 4, but not perfect, as known for a few decades (Herbig 1995). Nonetheless, the areas with high number density (in red) show smaller scatters and have better linear relationship (red solid lines). Even in sparsely sampled regions, such as those near low extinction of 0.15 mag and high extinction of over 0.6 mag, the medians do not deviate from the fitted curve. We do not include the samples with the higher extinction ($>$ 1 mag) in the fitting process due to the well-known “skin effect” (Herbig 1995), namely the strength of DIB cannot indefinitely increase with the extinction as expected. Besides, we also attempt a linear fit forced through the origin (green dashed lines) to facilitate the comparison with the previous works, for instance, Herbig (1975). It is clear that the green dashed lines are not suitable to represent the correlation of the DIBs with $\rm E(B-V)$, as they have larger fitting errors and deviate from the red solid lines.
Figure 10: Gaussian KDE number density plot of the EWs of the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ (from top to bottom) as a function of the extinction $\rm E(B-V)$ derived from Green et al. (2019). For each panel, the red solid line represents a linear fit to the median EW (which is marked with a white dot, and its standard deviation marked is with a black error bar) in each $\rm\Delta E(B-V)=0.05\ mag$ bin. The green dashed line is the same as the red solid line, but the intercept is forced to be zero. The coefficients of the linear fit and the Pearson correlation coefficient between the EW and $\rm E(B-V)$ are also annotated in each panel.
Table 4: Correlation between DIB and $\rm E(B-V)$ $(1)$$(1)$footnotemark: $(1)$